If two solutions are mixed with same kind of solute. How to find Solution 1 quantity in kg solvent if Solution 1 concentration has 3.975 m and solution 2 concentration has 6.532 m and solution 2 quantity has 4.937 in kg solvent and resulting to a mixed concentration of 5.421 m?
1
Expert's answer
2021-10-04T05:11:26-0400

Solution:

m1 = Molality of Solution 1 = 3.975 m

g1 = kg solvent = unknown

m2 = Molality of Solution 2 = 6.532 m

g2 = kg solvent = 4.937 kg

m = Molality of mixture = 5.421 m


m = (m1g1 + m2g2) / (g1 + g2)

5.421 = (3.975 × g1 + 6.532 × 4.937) / (g1 + 4.937)

5.421 = (3.975 × g1 + 32.2485) / (g1 + 4.937)

5.421 × (g1 + 4.937) = 3.975 × g1 + 32.2485

5.421 × g1 + 26.7635 = 3.975 × g1 + 32.2485

1.446 × g1 = 5.485

g1 = 3.793 kg


Answer: 3.793 kg